OpenStudy (anonymous):

The symmetry of the hyperbola (y-1)^2/9-(x-5)^2/16=1 occurs at the lines x=1 and y=5.

OpenStudy (anonymous):

does this hyperbola have an equation?

OpenStudy (anonymous):

The symmetry of the hyperbola (y-1)^2/9-(x-5)^2/16=1 occurs at the lines x=1 and y=5

OpenStudy (anonymous):

is there more to the question or is it a true and false?

OpenStudy (anonymous):

y-1=0 y=1 x-5=0 x=5 i think

OpenStudy (anonymous):

just true or false

OpenStudy (anonymous):

@v4xN0s

OpenStudy (anonymous):

its flase

OpenStudy (anonymous):

thank you!